OpenStudy (anonymous):

How do I differentiate v''t^2=0

OpenStudy (xapproachesinfinity):

what is that \[v"t^2=0\] differentiate with respect to what?

OpenStudy (anonymous):

The original question was this

OpenStudy (caozeyuan):

I learned second order ODE when i was in high school, but they dont teach reduction of order, so I am gonna take a guess

OpenStudy (caozeyuan):

Clearly, t is solvable from the given, so solve for t first. then plug back and solve for y. my guess is that you will get two values for y. One is the given t^2, the other will be your answer

OpenStudy (caozeyuan):

that's just my guess so no guarantee it works or it is right

OpenStudy (caozeyuan):

you mean how to solve for t or how to solve for y?

OpenStudy (anonymous):

the solution for y2

OpenStudy (amistre64):

https://www.youtube.com/watch?v=oQSFW8BIrY0

OpenStudy (anonymous):

in his example he uses v'' and v' in mine I have v'' and v and I'm not sure how to equal w to that

OpenStudy (amistre64):

t^2y'' -4ty' +6y = 0, y1 = t^2 let y2 = vt^2 y2' = v't^2 + 2vt y2'' = v''t^2 +4v't +2v plug them in ... t^2(t^2 v'' +4t v' +2v) -4t(t^2v' + 2t v) +6t^2 v = 0 t^4 v'' +4t^3 v' +2t^2 v -4t^3 v' -8t^2 v +6t^2 v = 0 combine like terms t^4 v'' +(4t^3 -4t^3) v' (-8t^2 +6t^2 +2t^2) v= 0

OpenStudy (amistre64):

so, t^4v'' = 0 is what i see this as

OpenStudy (anonymous):

yeah i was overthinking it when I redid it lol, but how do I continue from there

OpenStudy (amistre64):

solve for v of course

OpenStudy (caozeyuan):

坑爹啊!!!linear methods don't work on non-linear equations!!

OpenStudy (anonymous):

That's what I mean I don't know how to solve that.

OpenStudy (amistre64):

this is a linear equation ...

OpenStudy (caozeyuan):

OK, I guess I won't get banned using Chinese

OpenStudy (caozeyuan):

no, cuz they don't have constant coeffs

OpenStudy (amistre64):

lol, its all diamonds to me

OpenStudy (caozeyuan):

at least that's what my high school math teacher said

OpenStudy (amistre64):

this is a 'second-order' _linear_ ordinary differential equation

OpenStudy (amistre64):

are my vs primed correctly? chk the algebra to make sure its good to this point

OpenStudy (anonymous):

yeah it's correct I got that too

OpenStudy (anonymous):

I still don't know how the answer is t^3 :l

OpenStudy (amistre64):

reduce let w = v' w' = v'' t^4 w' = 0 let z = w' t^4z = 0 but thats seems redundant to me, but we can integrate both sides

OpenStudy (amistre64):

forget the z

OpenStudy (amistre64):

divide off the t^4 w' = 0 integrate, w = C for some arbitrary constant

OpenStudy (amistre64):

w = v' = C v' = C, integrate again v = t + c

OpenStudy (amistre64):

y2 = vt^2 y2 = (t+c)t^2 y2 = t^3+ct^2, drop the y1 parts of it y2 = t^3

OpenStudy (anonymous):

gottt it you're a life saver

OpenStudy (amistre64):

recall w' means dw/dt dw = 0 dt, and constants derive to 0 w = some constant ... etc

OpenStudy (anonymous):

yeah totally missed that when I was doing it thank youuu

OpenStudy (amistre64):

good luck :)